OpenStudy (anonymous):

If you have 330.0 mL of water at 25.00 °C and add 100.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (aaronq):

the thermal energy (heat) lost but the hot water is transferred to the cold water, (assuming that loss does not occur to any other object) we can say: \(-q_{hot}=q_{warm}\) from the calorimetry equation "\(q=m*c_p*\Delta T\)", we can get: \(-[m_1\Delta T]=m_2\Delta T\rightarrow -[m_1(T_f-T_i)]=m_2\Delta(T_f-T_i)\) because the heat capacities are equal (they are both liquid water), we can ignore it.

OpenStudy (aaronq):

the last part shouldn't have a \(\Delta\) sign. it should read: "\(-[m_1(T_f-T_i)]=m_2(T_f-T_i) \)"

OpenStudy (anonymous):

so wait, does that mean that finding water's specific heat is not necessary?

OpenStudy (aaronq):

no because they're both liquid water, so the \(C_p\) will just cancel out in the math.

OpenStudy (anonymous):

so by inserting the data, the equation would be -(330.0(T-25))=100(t-95)?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

and then just solve for T to find the final temperature?

OpenStudy (aaronq):

yep